The machining division of ITA International has a capacity of 2,000 units. Its sales and cost data are:

Selling price per unit $80
Variable manufacturing costs per unit 25
Variable selling costs per unit 3
Total fixed manufacturing overhead 183,200



The machining division is currently selling 1,800 units to outside customers, and the assembly division of ITA International wants to purchase 400 units from machining. If the transaction takes place, the variable selling costs per unit on the units transferred to assembly will be $0/unit, and not $3/unit. If the assembly division is currently buying from an outside supplier at $75 per unit, what will be the effect on overall company profits if internal sales for 400 units take place at the optimum transfer price?

The company profits would increase or decrease by $( )
Answer #1

Spare capacity = 2000 -1800 = 200

Contribution per unit = 80 -25 -3 = 52

Transfer price =( Variable cost + Lost contribution ) / No of units to be sold

= [ ( 400 *25) + ( 200* 52) ] / 400 = 51

Overall increase in profits = $0

External Internal and external
Sales 144000 148400
Variable costs
Manufacturing 45000 50000
Selling 5400 4800
Fixed overheads 183200 183200
Profit / loss -89600 -89600